### v1

# OSMa-Bench: Evaluating Open Semantic Mapping Under Varying Lighting Conditions

Abstract
Open Semantic Mapping (OSM) is a key technology in robotic perception, combining semantic segmentation and SLAM techniques. This paper introduces a dynamically configurable and highly automated LLM/LVLM-powered pipeline for evaluating OSM solutions called OSMa-Bench (Open Semantic Mapping Benchmark). The study focuses on evaluating state-of-the-art semantic mapping algorithms under varying indoor lighting conditions, a critical challenge in indoor environments. We introduce a novel dataset with simulated RGB-D sequences and ground truth 3D reconstructions, facilitating the rigorous analysis of mapping performance across different lighting conditions. Through experiments on leading models such as ConceptGraphs [gu2024conceptgraphs], BBQ [linok2024beyond] and OpenScene [peng2023openscene], we evaluate the semantic fidelity of object recognition and segmentation. Additionally, we introduce a Scene Graph evaluation method to analyze the ability of models to interpret semantic structure. The results provide insights into the robustness of these models, forming future research directions for developing resilient and adaptable robotic systems. Our code is available at https://be2rlab.github.io/OSMa-Bench/.
I Introduction
Open-vocabulary semantic mapping (OSM) is a rapidly advancing field that lies at the intersection of robotics, computer vision, and natural language processing. OSM enables robots to construct rich, semantically meaningful representations of their environment, empowering them to navigate, and interact in more efficient manner. Recent state-of-the-art approaches have made impressive strides in leveraging vision-language models (VLMs) and novel 3D scene reconstruction techniques to generate highly expressive and interpretable spatial representations, making robots for general purposes one step closer [chang2023context, ding2023pla, gu2024conceptgraphs, huang2023openins3d, takmaz2023openmask3d, peng2023openscene]. However, several challenges remain to be addressed to fully utilize OSM in real-world robotics applications. These include maintaining consistency in local updates, improving the clarity of semantic boundaries, adapting models to specific domains, increasing computational efficiency, and resolving ambiguities in natural language. Critically, many VLMs are trained on large-scale web datasets of natural images, yet robotic environments often exhibit significantly different visual characteristics. Indoor settings, in particular, can exhibit dramatic variations in illumination throughout the day and from artificial lighting. Domain-specific visual artifacts introduce additional complications. For OSM systems to be practical, they must be able to handle these diverse and dynamic conditions robustly.
Figure 1: Our work evaluates open semantic mapping quality providing an automated LLM/LVLM-based alternative to human assessment. We generate test sequences with different lighting conditions in a simulated indoor environment. Extra modifier such as variations in robot nominal velocity is applied as well.
II Related Work
Recent advances in open-vocabulary semantic mapping have leveraged VLMs and multimodal fusion techniques to enable robots to interpret unstructured environments using free-form language queries. Existing approaches can be broadly categorized by their technical paradigms, evaluation methodologies, and robustness considerations, which we analyze below to distinguish our benchmarking framework.
Open-Vocabulary 3D Scene Understanding
Contemporary methods typically employ 2D-3D feature fusion strategies to transfer open-vocabulary capabilities from vision-language models into 3D representations. ConceptFusion [jatavallabhula2023conceptfusion] pioneered this paradigm by aggregating CLIP [radford2021learning] features from multiview RGB images into point clouds, enabling semantic queries on the reconstructed scenes. While their evaluation on the UnCoCo dataset demonstrated multimodal query capabilities (text, image, audio), robustness under varying environmental conditions was not systematically assessed. ConceptGraphs [gu2024conceptgraphs] extended this approach by constructing hierarchical 3D scene graphs that encode spatial relationships and affordances, evaluated through human studies on Amazon Mechanical Turk. Similarly, OpenFusion [yamazaki2024open] optimized real-time performance through efficient feature fusion, prioritizing frame rates (FPS metrics) over illumination robustness. These works primarily benchmark on ScanNet [dai2017scannet], Replica [straub2019replica], and AI2-THOR [kolve2017ai2] datasets—environments captured under controlled lighting conditions that do not test operational reliability in dynamic settings.
Language-Guided 3D Instance Segmentation
Several methods focus on grounding language queries in 3D geometry. OpenScene [peng2023openscene] co-embeds 3D points with CLIP features for zero-shot segmentation, achieving mIoU scores of 36.1% on ScanNet and 51.2% on Matterport3D through cross-modal alignment. OpenMask3D [takmaz2023openmask3d] combines class-agnostic 3D masks with 2D open-vocabulary classifiers, attaining 10.7% AP50 on ScanNet200’s long-tail classes. While these methods demonstrate open-vocabulary capabilities, their evaluations assume consistent illumination and sensor characteristics—conditions rarely met in real-world deployments.
Robustness-Oriented Semantic Mapping
Recent work has begun addressing environmental variability through specialized architectures. BBQ [linok2024beyond] integrates LLMs with 3D scene graphs to resolve ambiguous queries in multi-instance environments, though computational efficiency remains a limitation. PLA [ding2023pla] enhances geometric consistency through planar constraints, improving mAP50 by 4.2% over baselines on ScanNet and S3DIS [armeni20163d]. Semantic Abstraction [ha2022semantic] employs LLM-driven relevance filtering to reduce voxel-level prediction noise in AI2-THOR environments. However, these approaches still evaluate under idealized lighting conditions, leaving their performance in low-light or dynamically lit scenarios unverified. GQA (Graph Question Answering) [hudson2019gqa] introduces a structured dataset for real-world visual reasoning, leveraging scene graphs from the Visual Genome dataset [krishna2017visual]. It covers object identification, attributes, relationships, spatial localization, and logical reasoning, ensuring consistency and interpretability in question-answering tasks. Space3D-Bench [szymanska2024space3d] extends 3D QA to spatial domains, evaluating models on object location, measurement, navigation, and scene patterns. It employs a two-tier evaluation: Ground Truth Check for fixed-answer questions and Answer Cross-Check using LLM-based validation for subjective responses.
Evaluation methodologies of current benchmarks focus narrowly on segmentation accuracy (mIoU, mAP50) and geometric fidelity, neglecting operational factors critical for robotics. For instance, OpenIns3D [huang2023openins3d] evaluates instance segmentation across five datasets using AP25/AP50 metrics but does not quantify performance degradation under sensor noise or motion blur. OVSG [chang2023context] introduces grounding success rate metrics for spatial queries but assumes static illumination during testing. The absence of standardized robustness evaluation protocols—particularly for lighting variations and sensor dynamics—creates a critical gap in assessing real-world viability.
Among the various skills assessed by current benchmarking methods, our current focus is on evaluating four key areas: 3D semantic segmentation, object existence determination, detection completeness, and spatial reasoning.
Figure 2: Evaluation Pipeline Diagram. We use Habitat Sim to establish test configurations that include various environmental factors like lighting and agent trajectories, generating a diverse dataset. This allows for tailored scenes, with randomly initialized agent trajectories simulating realistic interactions for effective model evaluation.
III Main Contribution
Figure 3: Replica CAD Dataset with Augmented Semantics. We expanded the semantic description of the replica CAD dataset. This made it possible to take into account during testing both classes describing parts of the apartment (for example, “wall”, “floor”, “stairs”) and classes describing furniture and household utensils.
OSMa-Bench (Open Semantic Mapping Benchmark) introduces a novel framework for evaluating robustness in open semantic mapping, addressing gaps in indoor environment simulation and multimodal reasoning assessment. Our key contributions are:
  1. 1.
    Dynamic Lighting Datasets with Augmented Ground Truth. We extend two datasets with controlled perturbations:
    • ReplicaCAD [szot2021habitat]. Nominal illumination with other three lighting variants (baseline, dynamic, camera light). We augmented per-instance masking GT to resolve occlusion ambiguities (Fig. 3).
    • HM3D [ramakrishnan2021habitat]. Nominal illumination with another variant - camera light conditions.
    Both combined with variable agent velocities, which enables systematic analysis of illumination/speed-induced performance degradation.
  2. 2.
    Automated Knowledge Graph Benchmarking. We develop an LLM-driven pipeline using Google Gemini [team2024gemini] to:
    • generate context-aware questions (via in-context templates) tailored to graph structures;
    • validate answers using hybrid metrics: binary (confusion matrices) for yes/no queries, and LLM-based consistency checks for open-ended attributes.
  3. 3.
    Cross-Method Robustness Analysis. We benchmark three SOTA methods [peng2023openscene, gu2024conceptgraphs, linok2024beyond] showing the effect of case conditions on their accuracy and reasoning capabilities.
OSMa-Bench unifies dynamic simulation, automated semantic validation, and reproducible evaluation, providing actionable insights for developing robust OSM systems.
Compared to many existing benchmarks that use static validation datasets, the proposed methodology reduces the risk of overfitting and data contamination for fair comparisons between methods.This is achieved through dynamically configured test sequences that adjust scene conditions and context-aware set of queries for checking knowledge graph quality.
Relying on LLM/LVLM capabilities and thanks to the suggested approach in general the benchmark achieves a high level of automation. This enables extensive large-scale testing, allows incremental extension of OSM skills to be tested, facilitates adaptation to various environments and target use-cases, all while keeping low maintenance costs.
IV Methodology
It is crucial to have standardized datasets for testing and evaluating different models and algorithms. However, due to the diversity of real-world scenes and scenarios, creating a comprehensive dataset that covers all possible variations and challenges is a complex task. To address this challenge, we designed an automated LLM/LVLM benchmarking pipeline (see Fig. 2). Our proposed evaluation methodology encompasses multiple facets to thoroughly assess the performance of open semantic mapping methods.
One of the ideas of this work is to analyze sensitivity of SOTA approaches to conditions modifications and identify what methods are most significantly impacted by these changes in terms of semantic segmentation and knowledge graph quality.
In particular, we introduce an approach for synthetically reproducing lighting conditions in simulation environment, e.g. placing the primary light source on the camera, introducing dynamic lighting varying along robot’s path, and removing external light sources. Tweaking these parameters allows for the creation of a diverse set of customized scenes tailored to specific testing scenarios. For example, different times of day can be simulated by adjusting the lighting settings.
Additionally, we vary robotic agent reference velocities and automatically generate movement trajectories within the scene, with random initialization of start and goal poses. This allows for the simulation of realistic agent behaviors and interactions with the environment and provides a more dynamic and challenging testing environment.
Each semantic map is compared against ground truth data, with the baseline configuration serving as the reference performance point. Deviations from this baseline highlight the sensitivity of methods to environmental changes.
IV-A Sematic Segmentation Evaluation
Figure 4: VQA Pipeline Diagram. We utilize LLM and LVLM to get scene frames descriptions and construct a set of questions and corresponding ground truth answers, which are additionally validated and balanced in order to avoid usage of ambiguous questions for the further evaluation of the scene graph.
Segmentation evaluation aims to assess the quality of segmentation of 3D scenes using a open set of classes present in our datasets. Here we employ the mean accuracy (mAcc) and frequency-weighted mean Intersection over Union (f-mIoU) metrics.
mAcc represents the average accuracy across all classes
$mAcc=\frac{1}{C}\sum_{c=1}^{C}\frac{TP_{c}}{TP_{c}+FN_{c}}$ (1)
where $TP_{c}$ and $FP_{c}$ are the number of true positive and false positive predictions for class $c$, respectively.
f-mIoU provides a weighted measure, taking into account both the per-class Intersection over Union (IoU) and the frequency of each class in the dataset.
$f\text{-}mIoU=\frac{\sum_{c=1}^{C}n_{c}\cdot IoU_{c}}{\sum_{c=1}^{C}n_{c}}$ (2)
where $C$ is the total number of classes, $n_{c}$ is the number of points for class $c$, and $IoU_{c}$ is the Intersection over Union (IoU) for class $c$.
While f-mIoU is a commonly preferred metric for semantic segmentation tasks due to its ability to handle imbalanced classes effectively, there may be scenarios where mAcc shows the quality of segmentation masks. The set of suggested metrics can be also extended to include almost any value based on confusion matrix.
IV-B Semantic Graph Evaluation
To assess the consistency and accuracy of semantic graphs we developed the question-answer generation pipeline, presented in Figure 4. This pipeline is functioning in the following way.
We sample frames along the movement within a previously generated test scene. For each key frame, we employ LVLM to generate scene descriptions and further construct a set of questions each targeting specific aspects of scene understanding:
  1. 1.
    Binary General – Yes/No questions about the presence of objects and general scene characteristics (e.g., “Is there a blue sofa?”).
  2. 2.
    Binary Existence-Based – Yes/No questions designed to track false positives by querying non-existent objects (e.g., “Is there a piano?”).
  3. 3.
    Binary Logical – Yes/No questions with logical operators such as AND/OR (e.g., “Is there a chair AND a table?”).
  4. 4.
    Measurement – Questions requiring numerical answers related to object counts or scene attributes (e.g., “How many windows are present?”).
  5. 5.
    Object Attributes – Queries about object properties, including color, shape, and material (e.g., “What color is the door?”).
  6. 6.
    Object Relations - Functional – Questions about functional relationships between objects (e.g., “Which object supports the table?”).
  7. 7.
    Object Relations - Spatial – Queries about spatial placement of objects within the scene (e.g., “What is in front of the staircase?”).
  8. 8.
    Comparison – Questions that compare object properties such as size, color, and position (e.g., “Which is taller: the bookshelf or the lamp?”).
Next, LVLM-generated descriptions are aggregated into a unified scene description followed by a validation step for QA sets using a combination of LVLM and LLM to detect inconsistencies. Duplicate questions are filtered and answers ambiguities gets resolved, ensuring logical consistency across frames. A balancing mechanism regulates objects frequency to prevent over-representation. The final validated question set is stored for evaluation.
We evaluate answer correctness using the Answering Accuracy metric. For binary and measurement-based questions, answers are directly compared to the ground truth. Attribute- and search-related questions’ correctness is determined through the LLM-based semantic evaluation. LLM assesses whether the predicted answer aligns with the ground truth, returning a structured JSON output. The final Accuracy is computed as the ratio of correct answers to the total number of questions.
We performed prompt engineering to design question generation, validation, and evaluation through LLM and LVLM. Structured prompts were developed for scene description, ensuring consistent capture of object attributes and spatial relations; question generation, enforcing balanced category distribution and restricting inference beyond the provided scene description; and question validation, filtering out semantically invalid, ambiguous or logically inconsistent questions based on both the image and textual scene representation.
The suggested methodology allows to adapt tests conditions by changing agent reference trajectories, light conditions along the path, and other scene modifications as well as questions from the VQA pipeline for separate tests just by changing a few corresponding parameters in the configuration file.
V Experiments
V-A Experiments Organization
The first step in conducting experiments is the data preparation that involves loading 3D scene models, configuring lighting, and agent paths and velocity. The latter can done within Habitat Sim [szot2021habitat] by adjusting corresponding JSON files.
Two various datasets have been deployed for testing in Habitat-Sim environment [szot2021habitat]. The first one is a synthetic Replica CAD [szot2021habitat] dataset that was selected because of its flexibility in configuring test conditions and available precise ground truth for both object positions and scene semantics [straub2019replica]. The second one is a photo-realistic Habitat-Matterport 3D Research Dataset (HM3D) [ramakrishnan2021habitat], which is one of the largest collection of high-resolution indoor 3D scans.
We assigned sets of 22 and 8 scenes for ReplicaCAD [szot2021habitat] and HM3D [ramakrishnan2021habitat] datsets respectively distinct in test conditions’ configurations:
  • pairs of start and goal robot poses have been assigned randomly from an available navigation mesh to cover most of the available scene;
  • collision-free path between terminal states have been defined relying on a GreedyGeodesicFollower motion planner available in Habitat-Sim [ramakrishnan2021habitat];
  • nominal value of robot’s constant reference velocity along a path have been assigned to 0.75 m/s and 0.8 rad/s, while for a ’velocity’ modification we doubled these values;
  • lighting conditions have been set to the following scenarios:
    • ”baseline” configuration uses the static not-uniformly distributed light sources available as a default scenario for the Replica CAD dataset only;
    • ”nominal lights” configuration relies on the mesh itself emitting light without any added light sources;
    • ”camera light” introduced extra directed light source attached to the camera;
    • ”dynamic lighting” corresponds to changing light conditions along robot path.
To implement LLM and VLM blocks within the VQA pipeline depicted on Fig. 4 we used Google Gemini 2.0-Flash [team2024gemini].
The VQA pipeline generated on average 184 and 76 questions for each scene of ReplicaCAD [szot2021habitat] and HM3D [ramakrishnan2021habitat] respectively, which gives us in total 4055 and 611 questions for ReplicaCAD [szot2021habitat] and HM3D [ramakrishnan2021habitat] respectively. The average ratio between different questions’ categories is the the following: 18.6% for Binary General, 16.6% for Binary Existence-Based, 18.4% for Binary Logical, 5.2% for Measurement, 17.0% for Object Attributes, 0.8% for Object Relations - Functional, 18.7% for Object Relations - Spatial, 4.7% for Comparison.
Functional Relationships were challenging for LLM to interpret correctly, often leading to inconsistent or ambiguous answers. As a result, many of these questions were removed during the validation process, leaving only a small proportion in the final datasets. Due to the low number of these questions, their inclusion would make comparisons with other categories unreliable. Thus, we exclude them from evaluation.
For more fair comparison between different methods we replaced ChatGPT4 LLM originally suggested by ConceptGraphs [gu2024conceptgraphs] by LLaMa3.1-8B [dubey2024llama] also used by BBQ [linok2024beyond]. While LLM plays an important role in building a semantic graph, it is not a part of reported contributions, and most approaches use it of-the-shelf. Moreover, we wanted to verify performance given more compact model.
V-B Experimental Results
The evaluation was conducted for three methods: ConceptGraphs [gu2024conceptgraphs], OpenScene [peng2023openscene], and a very recent BBQ [linok2024beyond]. These methods employ different approaches, so we can perform benchmarking on a wider spectrum.
Obtained results for segmentation quality metrics for all considered methods and scenes are presented in Tables I-III, while its relative change under different test conditions are illustrated on a ReplicaCAD [szot2021habitat] dataset by Fig. 5.
Semantic graph quality assessment have been performed according to the described methodology. Fig 6 illustrates VQA accuracy achieved by BBQ [linok2024beyond] and ConceptGraphs [gu2024conceptgraphs] methods for different categories of semantic queries.
For BBQ [linok2024beyond], question related to object relations were not considered, as its scene graph construction is query-dependent. Instead, the evaluation focused on assessing the correctness of object presence and attributes, ensuring that the retrieved object set aligns with the given question.
TABLE I: mAcc Values on ReplicaCAD($\uparrow$)
Approach FPS (Hz) Baseline Camera Light Dynamic Lights Nominal Lights Velocity Min Max Avg
ConceptGraphs[gu2024conceptgraphs] $0.19^{*}$ 0.291 0.276 0.260 0.258 0.283 0.258 0.291 0.274
BBQ[linok2024beyond] $0.93^{*}$ 0.162 0.165 0.128 0.157 0.140 0.128 0.165 0.150
Openscene[peng2023openscene] 0.13 0.130 0.127 0.140 0.136 0.132 0.127 0.140 0.133
- including scene graph generation
TABLE II: f-mIoU Values on ReplicaCAD ($\uparrow$)
Approach Baseline Camera Light Dynamic Lights Nominal Lights Velocity Min Max Avg
ConceptGraphs[gu2024conceptgraphs] 0.324 0.296 0.299 0.310 0.309 0.296 0.324 0.308
BBQ[linok2024beyond] 0.238 0.232 0.179 0.205 0.213 0.179 0.238 0.213
Openscene[peng2023openscene] 0.484 0.482 0.479 0.469 0.475 0.469 0.484 0.478
TABLE III: Semantic Segmentation on HM3D
Approach FPS (Hz) Camera Light Nominal Lights Velocity
mAcc $\uparrow$ f-mIoU $\uparrow$ mAcc $\uparrow$ f-mIoU $\uparrow$ mAcc $\uparrow$ f-mIoU $\uparrow$
ConceptGraphs[gu2024conceptgraphs] $0.16^{*}$ 0.143 0.150 0.158 0.169 0.156 0.156
BBQ[linok2024beyond] $0.92^{*}$ 0.085 0.202 0.124 0.207 0.087 0.187
Openscene[peng2023openscene] 0.25 0.116 0.293 0.118 0.352 0.121 0.344
- including scene graph generation
Figure 5: Comparison of degradation for mAcc and f-mIoU across different environmental conditions (less means less robust). The degradation is computed as $D(\%)=\frac{C-B}{B}\times 100$, where $B$ is the baseline metric and $C$ is the metric under the given condition.
Figure 6: Comparison of Answering Accuracy. We evaluate BBQ  [linok2024beyond] and ConceptGraphs [gu2024conceptgraphs] across ReplicaCAD [szot2021habitat] and HM3D [ramakrishnan2021habitat] sequences for different configurations.
In addition, we measure time consumption using NVIDIA GeForce RTX 4090 for 3D scene representation in frames per second (FPS), which is shown in Table  I and Table  III.
V-C Results Analysis
Through this systematic variation and evaluation, we aimed to comprehensively understand the robustness and reliability of SOTA methods under different environmental conditions and configurations. The major results are:
  • ConceptGraphs [gu2024conceptgraphs] showed the best value for the mAcc metric on all datasets, but still struggles to accurately detect small objects.
  • BBQ [linok2024beyond], taking into account the resulted FPS, seemed to be faster than ConceptGraphs [gu2024conceptgraphs], while BBQ [linok2024beyond] has a significant decrease in metric values under different lighting conditions and turned out to be the most sensitive.
  • OpenScene [peng2023openscene] turned out to be the most stable method among those tested (best f-mIoU overall). The advantage of this approach is a rich semantic description of the scene. However, the quality of the constructed semantic masks is quite poor. We assume that it happens due to the usage of pixel-aligned embeddings, not a segmentation model.
Figure 6 presents the answering accuracy across different question categories and scene modifications. Binary existence-based questions exhibit high accuracy, indicating that models do not hallucinate non-existent objects. Overall, binary questions achieve accuracy above 50–60%, confirming non-random response patterns.
Figure 7: Illustration of common corner cases in 3D semantic mapping under varying environmental conditions. Issues include segmentation errors due to illumination changes , loss of small objects during reconstruction, noisy point clouds, misclassification of shadows as objects, and incorrect object merging leading to inflated instance counts in scene graphs.
The highest accuracy is observed under nominal conditions — the baseline for ReplicaCAD [szot2021habitat] and nominal lighting for HM3D [ramakrishnan2021habitat]. Variations in lighting and velocity reduce accuracy, especially in object attributes and relationships questions. Attribute-based questions often fail due to missing material and property details in scene graph captions, leading to incomplete LLM responses.
Taking into account the analysis of metrics, it was also possible to detect the following corner cases under variable lighting conditions, which can lead to a decrease in accuracy. Examples of the following corner cases are shown in Fig.  7. When using camera light, shadows of objects can be counted as part of the object. Dynamic lighting causes incorrect segmentation of uniform surfaces, such as floors, due to illumination differences. Also, when using only the camera light source, a noisy point cloud was obtained. The tested methods show sensitivity to objects with a standard description (for example, “wall cabinet” and “base cabinet”). Also, when assessing the quality of semantic segmentation, it turned out that the methods show worse results if we use a ‘_’ instead of a space in the class name.
Object merging issues occur, for example in ConceptGraphs on ReplicaCAD, where staircases are split into individual steps. This impacts Measurement-based question answering, e.g.:
Question: How many staircases are present?
Answer: 1
Category: Measurement
Scene Graph Answer: 15
Similar errors appear in BBQ, inflating instance counts.
VI Conclusion
In this paper, we target a systematic evaluation of open semantic mapping algorithms, in particular their robustness in dynamic environments, e.g. under varying indoor lighting conditions. To accomplish the task we introduced a dynamically configurable and highly automated LLM/LVLM-powered benchmarking methodology.
Our novel dataset enabled a comprehensive analysis of leading methods, such as OpenScene [peng2023openscene], BBQ [linok2024beyond], and Concept Graph [gu2024conceptgraphs], highlighting the impact of lighting variations on semantic segmentation accuracy and quality of constructed knowledge graphs.
Analysis of the obtained results and corner cases points directions for further research in developing more resilient algorithms capable of maintaining high performance across diverse environments.
In our future work, we aim to expand the proposed methodology to assess a broader range of OSM methods focusing on skills crucial for autonomous navigation and general-purpose robots in dynamic environments. In particular, benchmarking visual grounding and scene graph evaluation skills will require adapting the suggested VQA pipeline either to generate question based on ground truth info on all instances and its positions (if available from the dataset) or to process a complete scene representation rather than analyzing it frame-by-frame. Verifying actionable affordances, spatial reasoning, and traversability skills will require only extending questions’ categories within the VQA pipeline, i.e. adapting the service prompts. We also did a preliminary study on incorporating synthetic dynamic objects into scenes and will continue to develop this aspect for the upcoming edition of OSMa-Bench.
[…]

### v2

@@
 We introduce a novel dataset with simulated RGB-D sequences and ground truth 3D reconstructions, facilitating the rigorous analysis of mapping performance across different lighting conditions.
-Through experiments on leading models such as ConceptGraphs [gu2024conceptgraphs], BBQ [linok2024beyond] and OpenScene [peng2023openscene], we evaluate the semantic fidelity of object recognition and segmentation.
+Through experiments on leading models such as ConceptGraphs [1], BBQ [2], and OpenScene [3], we evaluate the semantic fidelity of object recognition and segmentation.
 Additionally, we introduce a scene graph evaluation method to analyze the ability of models to interpret semantic structure.
@@
 I Introduction
-Open-vocabulary semantic mapping (OSM) is a rapidly advancing field that lies at the intersection of robotics, computer vision, and natural language processing.
-OSM enables robots to construct rich, semantically meaningful representations of their environment, empowering them to navigate, and interact in more efficient manner.
-Recent state-of-the-art approaches have made impressive strides in leveraging vision-language models (VLMs) and novel 3D scene reconstruction techniques to generate highly expressive and interpretable spatial representations, making robots for general purposes one step closer [chang2023context, ding2023pla, gu2024conceptgraphs, huang2023openins3d, takmaz2023openmask3d, peng2023openscene].
+Open-vocabulary semantic mapping (OSM) is a rapidly advancing field in robotics, computer vision, and natural language processing.
+OSM enables robots to construct rich, semantically meaningful representations of their environment, empowering them to navigate and interact more efficiently.
+Recent state-of-the-art approaches have made impressive strides in leveraging vision-language models (VLMs) and novel 3D scene reconstruction techniques to generate highly expressive and interpretable spatial representations, making robots for general purposes one step closer [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13].
 However, several challenges remain to be addressed to fully utilize OSM in real-world robotics applications.
@@
 Critically, many VLMs are trained on large-scale web datasets of natural images, yet robotic environments often exhibit significantly different visual characteristics.
-Indoor settings, in particular, can exhibit dramatic variations in illumination throughout the day and from artificial lighting.
+Indoor settings can experience dramatic variations in illumination throughout the day and from artificial lighting.
 Domain-specific visual artifacts introduce additional complications.
 For OSM systems to be practical, they must be able to handle these diverse and dynamic conditions robustly.
-Figure 1: Our work evaluates open semantic mapping quality providing an automated LLM/LVLM-based alternative to human assessment.
+Figure 1: Our work evaluates open semantic mapping quality, providing an automated LLM/LVLM-based alternative to human assessment.
 We generate test sequences with different lighting conditions in a simulated indoor environment.
-Extra modifier such as variations in robot nominal velocity is applied as well.
+Extra modifier, such as variations in robot nominal velocity, is applied as well.
 II Related Work
-Recent advances in open-vocabulary semantic mapping have leveraged VLMs and multimodal fusion techniques to enable robots to interpret unstructured environments using free-form language queries.
-Existing approaches can be broadly categorized by their technical paradigms, evaluation methodologies, and robustness considerations, which we analyze below to distinguish our benchmarking framework.
+Recent progress in open-vocabulary semantic mapping leverages VLMs and multimodal fusion to enable robots to interpret scenes via natural language.
+Existing approaches fall into three categories: 3D scene understanding, language-guided segmentation, and robustness-oriented mapping.
 Open-Vocabulary 3D Scene Understanding
-Contemporary methods typically employ 2D-3D feature fusion strategies to transfer open-vocabulary capabilities from vision-language models into 3D representations.
-ConceptFusion [jatavallabhula2023conceptfusion] pioneered this paradigm by aggregating CLIP [radford2021learning] features from multiview RGB images into point clouds, enabling semantic queries on the reconstructed scenes.
-While their evaluation on the UnCoCo dataset demonstrated multimodal query capabilities (text, image, audio), robustness under varying environmental conditions was not systematically assessed.
-ConceptGraphs [gu2024conceptgraphs] extended this approach by constructing hierarchical 3D scene graphs that encode spatial relationships and affordances, evaluated through human studies on Amazon Mechanical Turk.
-Similarly, OpenFusion [yamazaki2024open] optimized real-time performance through efficient feature fusion, prioritizing frame rates (FPS metrics) over illumination robustness.
-These works primarily benchmark on ScanNet [dai2017scannet], Replica [straub2019replica], and AI2-THOR [kolve2017ai2] datasets—environments captured under controlled lighting conditions that do not test operational reliability in dynamic settings.
+Most methods adopt 2D-3D feature fusion to project CLIP [14] features in 3D.
+ConceptFusion [4] aggregates image features into point clouds to enable semantic queries, while ConceptGraphs [1] and HOV-SG [5] extend this by constructing hierarchical scene graphs that capture object relations.
+OpenFusion [6] improves runtime efficiency, but emphasizes FPS over robustness.
+These methods are commonly benchmarked on ScanNet [15] or Replica [16] datasets with controlled lighting, limiting generalization to dynamic environments.
 Language-Guided 3D Instance Segmentation
-Several methods focus on grounding language queries in 3D geometry.
-OpenScene [peng2023openscene] co-embeds 3D points with CLIP features for zero-shot segmentation, achieving mIoU scores of 36.1% on ScanNet and 51.2% on Matterport3D through cross-modal alignment.
-OpenMask3D [takmaz2023openmask3d] combines class-agnostic 3D masks with 2D open-vocabulary classifiers, attaining 10.7% AP50 on ScanNet200’s long-tail classes.
-While these methods demonstrate open-vocabulary capabilities, their evaluations assume consistent illumination and sensor characteristics—conditions rarely met in real-world deployments.
+Some approaches focus on grounding language in 3D geometry.
+OpenScene [3] co-embeds point features with CLIP for zero-shot segmentation, while OpenMask3D [7] and Open3DIS [8] fuse class-agnostic 3D masks with 2D classifiers.
+Despite strong performance, these models assume stable sensor and lighting conditions, which are rarely met in real-world applications.
 Robustness-Oriented Semantic Mapping
-Recent work has begun addressing environmental variability through specialized architectures.
-BBQ [linok2024beyond] integrates LLMs with 3D scene graphs to resolve ambiguous queries in multi-instance environments, though computational efficiency remains a limitation.
-PLA [ding2023pla] enhances geometric consistency through planar constraints, improving mAP50 by 4.2% over baselines on ScanNet and S3DIS [armeni20163d].
-Semantic Abstraction [ha2022semantic] employs LLM-driven relevance filtering to reduce voxel-level prediction noise in AI2-THOR environments.
-However, these approaches still evaluate under idealized lighting conditions, leaving their performance in low-light or dynamically lit scenarios unverified.
-GQA (Graph Question Answering) [hudson2019gqa] introduces a structured dataset for real-world visual reasoning, leveraging scene graphs from the Visual Genome dataset [krishna2017visual].
-It covers object identification, attributes, relationships, spatial localization, and logical reasoning, ensuring consistency and interpretability in question-answering tasks.
-Space3D-Bench [szymanska2024space3d] extends 3D QA to spatial domains, evaluating models on object location, measurement, navigation, and scene patterns.
-It employs a two-tier evaluation: Ground Truth Check for fixed-answer questions and Answer Cross-Check using LLM-based validation for subjective responses.
-Evaluation methodologies of current benchmarks focus narrowly on segmentation accuracy (mIoU, mAP50) and geometric fidelity, neglecting operational factors critical for robotics.
-For instance, OpenIns3D [huang2023openins3d] evaluates instance segmentation across five datasets using AP25/AP50 metrics but does not quantify performance degradation under sensor noise or motion blur.
-OVSG [chang2023context] introduces grounding success rate metrics for spatial queries but assumes static illumination during testing.
-The absence of standardized robustness evaluation protocols—particularly for lighting variations and sensor dynamics—creates a critical gap in assessing real-world viability.
-Among the various skills assessed by current benchmarking methods, our current focus is on evaluating four key areas: 3D semantic segmentation, object existence determination, detection completeness, and spatial reasoning.
+Many works in the literature target environmental variability.
+BBQ [2] uses LLMs with 3D scene graphs to resolve ambiguous queries, though with high compute costs.
+PLA [9] introduces planar constraints to enforce geometric consistency, and Semantic Abstraction [10] applies LLM-driven filtering to reduce noise.
+However, evaluations still lack low-light or dynamic condition testing.
+The evaluation methodologies of the current benchmarks focus narrowly on segmentation accuracy (mIoU, mAP50) and geometric fidelity, neglecting operational factors critical for robotics.
+For example, OpenIns3D [11] evaluates instance segmentation across five datasets using AP25/AP50 metrics but does not quantify performance degradation under sensor noise or motion blur.
+Existing validation of graph scenes representations rarely goes beyond grounding: OVSG [12] introduces a grounding success rate for spatial queries but assumes static illumination and lacks standardized robustness protocols for lighting variation and sensor dynamics; in practice, evaluations rely on manually curated datasets or human assessment.
+Outside mapping, structured 3D‑QA benchmarks such as GQA [17] and Space3D‑Bench [18] are commonly used, but still not in an automatic way.
+We address this gap by evaluating four critical robotic capabilities: semantic segmentation, object existence, detection completeness, and spatial reasoning - under conditions that simulate real-world uncertainty.
 Figure 2: Evaluation Pipeline Diagram.
@@
 This made it possible to take into account during testing both classes describing parts of the apartment (for example, “wall”, “floor”, “stairs”) and classes describing furniture and household utensils.
-OSMa-Bench (Open Semantic Mapping Benchmark) introduces a novel framework for evaluating robustness in open semantic mapping, addressing gaps in indoor environment simulation and multimodal reasoning assessment.
+OSMa-Bench (Open Semantic Mapping Benchmark) introduces a novel framework for evaluating robustness in open semantic mapping, addressing gaps in an indoor environment simulation and multimodal reasoning assessment.
 Our key contributions are:
   1. 1.
-    Dynamic Lighting Datasets with Augmented Ground Truth. We extend two datasets with controlled perturbations:
-    • ReplicaCAD [szot2021habitat].
-Nominal illumination with other three lighting variants (baseline, dynamic, camera light).
-We augmented per-instance masking GT to resolve occlusion ambiguities (Fig.
-3).
-    • HM3D [ramakrishnan2021habitat]. Nominal illumination with another variant - camera light conditions.
-    Both combined with variable agent velocities, which enables systematic analysis of illumination/speed-induced performance degradation.
+    OSMa-Bench Datasets: ReplicaCAD [19] and HM3D [20] extended with lighting modes and doubled/nominal robot speed, plus per-instance masks (Fig. 3).
   2. 2.
-    Automated Knowledge Graph Benchmarking. We develop an LLM-driven pipeline using Google Gemini [team2024gemini] to:
-    • generate context-aware questions (via in-context templates) tailored to graph structures;
-    • validate answers using hybrid metrics: binary (confusion matrices) for yes/no queries, and LLM-based consistency checks for open-ended attributes.
+    LLM-based QA Pipeline: Gemini [21] generates and grades graph-aware questions, yielding fully automated semantic evaluation.
   3. 3.
-    Cross-Method Robustness Analysis.
-We benchmark three SOTA methods [peng2023openscene, gu2024conceptgraphs, linok2024beyond] showing the effect of case conditions on their accuracy and reasoning capabilities.
+    Cross-Method Robustness Analysis: We compare OpenScene, ConceptGraphs, and BBQ [3, 1, 2], showing the effect of case conditions on their accuracy and reasoning capabilities.
 OSMa-Bench unifies dynamic simulation, automated semantic validation, and reproducible evaluation, providing actionable insights for developing robust OSM systems.
-Compared to many existing benchmarks that use static validation datasets, the proposed methodology reduces the risk of overfitting and data contamination for fair comparisons between methods.This is achieved through dynamically configured test sequences that adjust scene conditions and context-aware set of queries for checking knowledge graph quality.
-Relying on LLM/LVLM capabilities and thanks to the suggested approach in general the benchmark achieves a high level of automation.
-This enables extensive large-scale testing, allows incremental extension of OSM skills to be tested, facilitates adaptation to various environments and target use-cases, all while keeping low maintenance costs.
+Unlike static benchmarks, OSMa-Bench harnesses LLM/LVLM automation to synthesize scene variants and context-aware queries on the fly, which both minimizes over-fitting and drives annotation cost toward zero.
+This high level of automation also makes the suite highly scalable: new lighting regimes or robot behaviours can be added with a few changes to the configuration, so even large, diverse experiments remain low-maintenance and domain-agnostic.
 IV Methodology
@@
 Our proposed evaluation methodology encompasses multiple facets to thoroughly assess the performance of open semantic mapping methods.
-One of the ideas of this work is to analyze sensitivity of SOTA approaches to conditions modifications and identify what methods are most significantly impacted by these changes in terms of semantic segmentation and knowledge graph quality.
-In particular, we introduce an approach for synthetically reproducing lighting conditions in simulation environment, e.g.
-placing the primary light source on the camera, introducing dynamic lighting varying along robot’s path, and removing external light sources.
+One of the key ideas of this work is to analyze the sensitivity of SOTA approaches to condition modifications and to identify which methods are most significantly impacted by these changes in terms of semantic segmentation and knowledge graph quality.
+In particular, we introduce an approach for synthetically reproducing lighting conditions in a simulation environment, such as placing the primary light source on the camera, introducing dynamic lighting that varies along the robot’s path, and removing external light sources.
 Tweaking these parameters allows for the creation of a diverse set of customized scenes tailored to specific testing scenarios.
@@
 Figure 4: VQA Pipeline Diagram.
-We utilize LLM and LVLM to get scene frames descriptions and construct a set of questions and corresponding ground truth answers, which are additionally validated and balanced in order to avoid usage of ambiguous questions for the further evaluation of the scene graph.
-Segmentation evaluation aims to assess the quality of segmentation of 3D scenes using a open set of classes present in our datasets.
-Here we employ the mean accuracy (mAcc) and frequency-weighted mean Intersection over Union (f-mIoU) metrics.
-mAcc represents the average accuracy across all classes
-$mAcc=\frac{1}{C}\sum_{c=1}^{C}\frac{TP_{c}}{TP_{c}+FN_{c}}$ (1)
-where $TP_{c}$ and $FP_{c}$ are the number of true positive and false positive predictions for class $c$, respectively.
-f-mIoU provides a weighted measure, taking into account both the per-class Intersection over Union (IoU) and the frequency of each class in the dataset.
-$f\text{-}mIoU=\frac{\sum_{c=1}^{C}n_{c}\cdot IoU_{c}}{\sum_{c=1}^{C}n_{c}}$ (2)
-where $C$ is the total number of classes, $n_{c}$ is the number of points for class $c$, and $IoU_{c}$ is the Intersection over Union (IoU) for class $c$.
-While f-mIoU is a commonly preferred metric for semantic segmentation tasks due to its ability to handle imbalanced classes effectively, there may be scenarios where mAcc shows the quality of segmentation masks.
-The set of suggested metrics can be also extended to include almost any value based on confusion matrix.
+We utilize LLM and LVLM to get scene frames descriptions and construct a set of questions and corresponding ground truth answers, which are additionally validated and balanced in order to avoid usage of ambiguous questions for further evaluation of the scene graph.
+We evaluate segmentation quality by assessing how well methods segment 3D scenes using our datasets generated using Habitat AI [19] simulator.
+Ground truth semantic point clouds are reconstructed using depth images, semantic images, and camera poses.
+Following the evaluation pipeline from ConceptGraphs [1], we establish a unique closed vocabulary for each scene to generate predictions and employ nearest neighbor matching between predicted and ground truth point clouds.
+The mAcc and f-mIoU metrics were chosen to evaluate the segmentation quality.
 IV-B Semantic Graph Evaluation
-To assess the consistency and accuracy of semantic graphs we developed the question-answer generation pipeline, presented in Figure 4. This pipeline is functioning in the following way.
+To assess the consistency and accuracy of semantic graphs, we developed the question-answer generation pipeline, presented in Figure 4.
 We sample frames along the movement within a previously generated test scene.
-For each key frame, we employ LVLM to generate scene descriptions and further construct a set of questions each targeting specific aspects of scene understanding:
+For each key frame, we employ LVLM to generate scene descriptions and further construct a set of questions, each targeting specific aspects of scene understanding:
   1. 1.
-    Binary General – Yes/No questions about the presence of objects and general scene characteristics (e.g., “Is there a blue sofa?”).
+    Binary General – Yes/No questions about the presence of objects and general scene characteristics (e.g., “Is there a blue sofa?”);
   2. 2.
-    Binary Existence-Based – Yes/No questions designed to track false positives by querying non-existent objects (e.g., “Is there a piano?”).
+    Binary Existence-Based – Yes/No questions designed to track false positives by querying non-existent objects (e.g., “Is there a piano?”);
   3. 3.
-    Binary Logical – Yes/No questions with logical operators such as AND/OR (e.g., “Is there a chair AND a table?”).
+    Binary Logical – Yes/No questions with logical operators such as AND/OR (e.g., “Is there a chair AND a table?”);
   4. 4.
-    Measurement – Questions requiring numerical answers related to object counts or scene attributes (e.g., “How many windows are present?”).
+    Measurement – questions requiring numerical answers related to object counts or scene attributes (e.g., “How many windows are present?”);
   5. 5.
-    Object Attributes – Queries about object properties, including color, shape, and material (e.g., “What color is the door?”).
+    Object Attributes – queries about object properties, including color, shape, and material (e.g., “What color is the door?”);
   6. 6.
-    Object Relations - Functional – Questions about functional relationships between objects (e.g., “Which object supports the table?”).
+    Object Relations (Functional) – questions about functional relationships between objects (e.g., “Which object supports the table?”);
   7. 7.
-    Object Relations - Spatial – Queries about spatial placement of objects within the scene (e.g., “What is in front of the staircase?”).
+    Object Relations (Spatial) – queries about spatial placement of objects within the scene (e.g., “What is in front of the staircase?”);
   8. 8.
     Comparison – questions that compare object properties such as size, color, and position (e.g., “Which is taller: the bookshelf or the lamp?”).
-Next, LVLM-generated descriptions are aggregated into a unified scene description followed by a validation step for QA sets using a combination of LVLM and LLM to detect inconsistencies.
-Duplicate questions are filtered and answers ambiguities gets resolved, ensuring logical consistency across frames.
-A balancing mechanism regulates objects frequency to prevent over-representation.
-The final validated question set is stored for evaluation.
+Scene captions from key-frames are merged, an LLM then filters duplicates, fixes ambiguities, and balances object frequencies before evaluation.
+Duplicate questions are filtered, and answer ambiguities are resolved, ensuring logical consistency across frames.
+A balancing mechanism prevents object over-representation.
 We evaluate answer correctness using the Answering Accuracy metric.
@@
 The final Accuracy is computed as the ratio of correct answers to the total number of questions.
-We performed prompt engineering to design question generation, validation, and evaluation through LLM and LVLM.
-Structured prompts were developed for scene description, ensuring consistent capture of object attributes and spatial relations; question generation, enforcing balanced category distribution and restricting inference beyond the provided scene description; and question validation, filtering out semantically invalid, ambiguous or logically inconsistent questions based on both the image and textual scene representation.
-The suggested methodology allows to adapt tests conditions by changing agent reference trajectories, light conditions along the path, and other scene modifications as well as questions from the VQA pipeline for separate tests just by changing a few corresponding parameters in the configuration file.
+Prompts enforce structured captions, balanced question types, and reject out-of-scope or ambiguous queries.
+The suggested methodology allows adapting test conditions by changing agent reference trajectories, light conditions along the path, and other modifications of a scene, as well as questions from the VQA pipeline.
 V Experiments
 V-A Experiments Organization
-The first step in conducting experiments is the data preparation that involves loading 3D scene models, configuring lighting, and agent paths and velocity.
-The latter can done within Habitat Sim [szot2021habitat] by adjusting corresponding JSON files.
-Two various datasets have been deployed for testing in Habitat-Sim environment [szot2021habitat].
-The first one is a synthetic Replica CAD [szot2021habitat] dataset that was selected because of its flexibility in configuring test conditions and available precise ground truth for both object positions and scene semantics [straub2019replica].
-The second one is a photo-realistic Habitat-Matterport 3D Research Dataset (HM3D) [ramakrishnan2021habitat], which is one of the largest collection of high-resolution indoor 3D scans.
-We assigned sets of 22 and 8 scenes for ReplicaCAD [szot2021habitat] and HM3D [ramakrishnan2021habitat] datsets respectively distinct in test conditions’ configurations:
-  • pairs of start and goal robot poses have been assigned randomly from an available navigation mesh to cover most of the available scene;
-  • collision-free path between terminal states have been defined relying on a GreedyGeodesicFollower motion planner available in Habitat-Sim [ramakrishnan2021habitat];
-  • nominal value of robot’s constant reference velocity along a path have been assigned to 0.75 m/s and 0.8 rad/s, while for a ’velocity’ modification we doubled these values;
+The first step in conducting experiments is data preparation, which involves loading 3D scene models, and configuring lighting, agent, and path.
+The latter can be done within Habitat Sim [19] by adjusting the corresponding JSON files.
+Two datasets have been deployed for testing in Habitat-Sim environment [19].
+The first one is a synthetic Replica CAD [19] dataset that was selected because of its flexibility in configuring test conditions and the availability of precise ground truth for both object positions and scene semantics.
+The second one is a photo-realistic Habitat-Matterport 3D Research Dataset (HM3D) [20], which is one of the largest collections of high-resolution indoor 3D scans.
+We assigned sets of 22 scenes from the ReplicaCAD dataset and 8 scenes from the HM3D dataset, each with distinct test condition configurations:
+  • Pairs of start and goal robot poses were randomly sampled from the available navigation mesh to ensure broad coverage of each scene;
+  • Collision-free path between terminal states has been defined relying on a GreedyGeodesicFollower motion planner available in Habitat-Sim;
   • Lighting conditions have been set to the following scenarios:
-    • ”baseline” configuration uses the static not-uniformly distributed light sources available as a default scenario for the Replica CAD dataset only;
-    • ”nominal lights” configuration relies on the mesh itself emitting light without any added light sources;
+    • ”baseline” configuration uses the static, not-uniformly distributed light sources available as a default scenario for the Replica CAD dataset only with baseline;
+    • ”nominal lights” configuration relies on the mesh itself emitting light without any added light sources. This means using baked lighting for HM3D and uniformly glowing meshes for ReplicaCAD;
     • ”camera light” introduced extra directed light source attached to the camera;
-    • ”dynamic lighting” corresponds to changing light conditions along robot path.
-To implement LLM and VLM blocks within the VQA pipeline depicted on Fig. 4 we used Google Gemini 2.0-Flash [team2024gemini].
-The VQA pipeline generated on average 184 and 76 questions for each scene of ReplicaCAD [szot2021habitat] and HM3D [ramakrishnan2021habitat] respectively, which gives us in total 4055 and 611 questions for ReplicaCAD [szot2021habitat] and HM3D [ramakrishnan2021habitat] respectively.
-The average ratio between different questions’ categories is the the following: 18.6% for Binary General, 16.6% for Binary Existence-Based, 18.4% for Binary Logical, 5.2% for Measurement, 17.0% for Object Attributes, 0.8% for Object Relations - Functional, 18.7% for Object Relations - Spatial, 4.7% for Comparison.
+    • ”dynamic lighting” corresponds to changing light conditions along robot path;
+  • Nominal value of robot’s constant reference velocities along path were set to 0.75 m/s for transition and 0.8 rad/s for rotation, while for a ”velocity” modification, we doubled these values for ReplicaCAD ”baseline” and HM3D ”nominal lights” conditions;
+To implement LLM/LVLM blocks within the VQA pipeline (Fig. 4), we used Google Gemini 2.0-Flash [21].
+The VQA pipeline generated on average 184 and 85 questions for each scene of ReplicaCAD and HM3D, which gives us in total 4055 and 681 questions for ReplicaCAD and HM3D, respectively.
+The average ratio between different questions’ categories is the following: 18.8% for Binary General, 16.4% for Binary Existence-Based, 18.1% for Binary Logical, 5.7% for Measurement, 16.9% for Object Attributes, 1.1% for Object Relations - Functional, 18.3% for Object Relations - Spatial, 4.8% for Comparison.
 Functional Relationships were challenging for LLM to interpret correctly, often leading to inconsistent or ambiguous answers.
-As a result, many of these questions were removed during the validation process, leaving only a small proportion in the final datasets.
-Due to the low number of these questions, their inclusion would make comparisons with other categories unreliable.
-Thus, we exclude them from evaluation.
-For more fair comparison between different methods we replaced ChatGPT4 LLM originally suggested by ConceptGraphs [gu2024conceptgraphs] by LLaMa3.1-8B [dubey2024llama] also used by BBQ [linok2024beyond].
-While LLM plays an important role in building a semantic graph, it is not a part of reported contributions, and most approaches use it of-the-shelf.
-Moreover, we wanted to verify performance given more compact model.
+As a result, many of these questions were removed during the validation process and, as a result, we excluded them from the evaluation.
 V-B Experimental Results
-The evaluation was conducted for three methods: ConceptGraphs [gu2024conceptgraphs], OpenScene [peng2023openscene], and a very recent BBQ [linok2024beyond].
+TABLE I: Performance on ReplicaCAD
+Approach FPS (Hz) Baseline Camera Light Dynamic Lights Nominal Lights Velocity
+mAcc f-mIoU mAcc f-mIoU mAcc f-mIoU mAcc f-mIoU mAcc f-mIoU
+ConceptGraphs [1] $0.19^{*}$ 29.53 28.00 29.76 22.71 24.73 14.29 26.80 21.34 28.32 27.33
+BBQ [2] $0.93^{*}$ 25.25 36.42 26.81 29.59 24.60 32.25 25.07 33.33 26.63 32.30
+OpenScene [3] 0.13 24.17 51.17 22.82 48.83 22.02 49.05 22.16 47.09 25.01 55.81
+- including scene graph generation
+TABLE II: Performance on HM3D
+Approach FPS (Hz) Nominal Lights Camera Light Velocity
+mAcc f-mIoU mAcc f-mIoU mAcc f-mIoU
+ConceptGraphs [1] $0.16^{*}$ 20.39 17.72 16.73 15.48 19.42 17.49
+BBQ [2] $0.92^{*}$ 18.14 23.96 17.24 24.03 18.28 24.71
+OpenScene [3] 0.25 20.58 34.19 17.29 29.97 19.77 34.02
+- including scene graph generation
+Figure 5: Comparison of changes (%) in mAcc and f-mIoU across conditions relative to the baseline on the ReplicaCAD dataset.
+Figure 6: Comparison of Answering Accuracy across ReplicaCAD and HM3D sequences.
+The evaluation was conducted for three methods: ConceptGraphs [1], OpenScene [3], and a very recent BBQ [2].
+We used 2D fused features for OpenScene, as this mode demonstrated the best performance.
 These methods employ different approaches, so we can perform benchmarking on a wider spectrum.
-Obtained results for segmentation quality metrics for all considered methods and scenes are presented in Tables I-III, while its relative change under different test conditions are illustrated on a ReplicaCAD [szot2021habitat] dataset by Fig.
+For more fair comparison between the methods, we replaced LLaVa-7b-v0 [22] and GPT4 [23] originally suggested by ConceptGraphs by LLaVa1.6-7b [24] and LLaMa3.1-8B [25], also used by BBQ.
+While LLM and LVLM play an important role in building a semantic graph, it is not a part of reported contributions, and most approaches use it off-the-shelf.
+Moreover, we wanted to verify performance given a more compact model.
+Obtained results for segmentation quality metrics for all considered methods and scenes are presented in Tables III, while their relative change under different test conditions is illustrated on a ReplicaCAD dataset by Fig.
 5.
-Semantic graph quality assessment have been performed according to the described methodology.
-Fig 6 illustrates VQA accuracy achieved by BBQ [linok2024beyond] and ConceptGraphs [gu2024conceptgraphs] methods for different categories of semantic queries.
-For BBQ [linok2024beyond], question related to object relations were not considered, as its scene graph construction is query-dependent.
+Final metrics are computed differently for each dataset: for ReplicaCAD, we calculate an overall confusion matrix since it provides a consistent closed vocabulary across all scenes, while for HM3D, we average metrics across individual scenes due to the large number of unique labels and potential label inconsistencies for identical objects across different scenes.
+Semantic graph quality assessment has been performed according to the described methodology.
+Fig 6 illustrates VQA accuracy achieved by BBQ and ConceptGraphs methods for different categories of semantic queries.
+We did not evaluate OpenScene because it does not construct a scene graph.
+For BBQ [2], questions related to object relations were not considered, as its scene graph construction is query-dependent.
 Instead, the evaluation focused on assessing the correctness of object presence and attributes, ensuring that the retrieved object set aligns with the given question.
-TABLE I: mAcc Values on ReplicaCAD($\uparrow$)
-Approach FPS (Hz) Baseline Camera Light Dynamic Lights Nominal Lights Velocity Min Max Avg
-ConceptGraphs[gu2024conceptgraphs] $0.19^{*}$ 0.291 0.276 0.260 0.258 0.283 0.258 0.291 0.274
-BBQ[linok2024beyond] $0.93^{*}$ 0.162 0.165 0.128 0.157 0.140 0.128 0.165 0.150
-Openscene[peng2023openscene] 0.13 0.130 0.127 0.140 0.136 0.132 0.127 0.140 0.133
-- including scene graph generation
-TABLE II: f-mIoU Values on ReplicaCAD ($\uparrow$)
-Approach Baseline Camera Light Dynamic Lights Nominal Lights Velocity Min Max Avg
-ConceptGraphs[gu2024conceptgraphs] 0.324 0.296 0.299 0.310 0.309 0.296 0.324 0.308
-BBQ[linok2024beyond] 0.238 0.232 0.179 0.205 0.213 0.179 0.238 0.213
-Openscene[peng2023openscene] 0.484 0.482 0.479 0.469 0.475 0.469 0.484 0.478
-TABLE III: Semantic Segmentation on HM3D
-Approach FPS (Hz) Camera Light Nominal Lights Velocity
-mAcc $\uparrow$ f-mIoU $\uparrow$ mAcc $\uparrow$ f-mIoU $\uparrow$ mAcc $\uparrow$ f-mIoU $\uparrow$
-ConceptGraphs[gu2024conceptgraphs] $0.16^{*}$ 0.143 0.150 0.158 0.169 0.156 0.156
-BBQ[linok2024beyond] $0.92^{*}$ 0.085 0.202 0.124 0.207 0.087 0.187
-Openscene[peng2023openscene] 0.25 0.116 0.293 0.118 0.352 0.121 0.344
-- including scene graph generation
-Figure 5: Comparison of degradation for mAcc and f-mIoU across different environmental conditions (less means less robust).
-The degradation is computed as $D(\%)=\frac{C-B}{B}\times 100$, where $B$ is the baseline metric and $C$ is the metric under the given condition.
-Figure 6: Comparison of Answering Accuracy.
-We evaluate BBQ  [linok2024beyond] and ConceptGraphs [gu2024conceptgraphs] across ReplicaCAD [szot2021habitat] and HM3D [ramakrishnan2021habitat] sequences for different configurations.
-In addition, we measure time consumption using NVIDIA GeForce RTX 4090 for 3D scene representation in frames per second (FPS), which is shown in Table  I and Table  III.
+In addition, we measure time consumption using NVIDIA GeForce RTX 4090 for 3D scene representation in frames per second (FPS), which is shown in Tables III.
 V-C Results Analysis
-Through this systematic variation and evaluation, we aimed to comprehensively understand the robustness and reliability of SOTA methods under different environmental conditions and configurations.
-The major results are:
-  • ConceptGraphs [gu2024conceptgraphs] showed the best value for the mAcc metric on all datasets, but still struggles to accurately detect small objects.
-  • BBQ [linok2024beyond], taking into account the resulted FPS, seemed to be faster than ConceptGraphs [gu2024conceptgraphs], while BBQ [linok2024beyond] has a significant decrease in metric values under different lighting conditions and turned out to be the most sensitive.
-  • OpenScene [peng2023openscene] turned out to be the most stable method among those tested (best f-mIoU overall).
-The advantage of this approach is a rich semantic description of the scene.
-However, the quality of the constructed semantic masks is quite poor.
-We assume that it happens due to the usage of pixel-aligned embeddings, not a segmentation model.
+Collecting mAcc and f-mIoU metrics from testing ConceptGraphs, BBQ, and OpenScene on scenes in ReplicaCAD and HM3D reveals key trade-offs.
+f-mIoU is weighted by class frequency, so voluminous objects like walls and floors matter more, while mAcc gives equal weight to all objects.
+On ReplicaCAD, OpenScene achieves the highest f-mIoU, performing best on voluminous objects.
+ConceptGraphs achieves the highest mAcc, indicating strength in recognizing more rare classes.
+BBQ performs steadily across both metrics but doesn’t lead in either.
+Under the velocity setting, OpenScene improves on both metrics, likely due to better filtering with fewer points.
+Under camera light, both BBQ and ConceptGraphs achieve higher scores in mAcc, likely because the direct lighting emphasizes objects, boosting detection.
+Also, ConceptGraphs is most sensitive to light conditions.
+On HM3D, performance drops across the board, but OpenScene shows the best performance.
 Figure 6 presents the answering accuracy across different question categories and scene modifications.
 Binary existence-based questions exhibit high accuracy, indicating that models do not hallucinate non-existent objects.
-Overall, binary questions achieve accuracy above 50–60%, confirming non-random response patterns.
+Overall, when binary questions achieve accuracy above 50–60%, it confirms non-random response patterns.
 Figure 7: Illustration of common corner cases in 3D semantic mapping under varying environmental conditions.
 Issues include segmentation errors due to illumination changes, loss of small objects during reconstruction, noisy point clouds, misclassification of shadows as objects, and incorrect object merging leading to inflated instance counts in scene graphs.
-The highest accuracy is observed under nominal conditions — the baseline for ReplicaCAD [szot2021habitat] and nominal lighting for HM3D [ramakrishnan2021habitat].
+The highest accuracy is observed under nominal conditions — the baseline for ReplicaCAD and nominal lighting for HM3D.
 Variations in lighting and velocity reduce accuracy, especially in object attributes and relationships questions.
 Attribute-based questions often fail due to missing material and property details in scene graph captions, leading to incomplete LLM responses.
-Taking into account the analysis of metrics, it was also possible to detect the following corner cases under variable lighting conditions, which can lead to a decrease in accuracy.
-Examples of the following corner cases are shown in Fig.
-7.
+Taking into account the analysis of metrics, it was also possible to detect corner cases (Fig.
+7) under variable lighting conditions, which can lead to a decrease in accuracy.
 When using camera light, shadows of objects can be counted as part of the object.
@@
 Also, when using only the camera light source, a noisy point cloud was obtained.
-The tested methods show sensitivity to objects with a standard description (for example, “wall cabinet” and “base cabinet”).
+The tested methods show sensitivity to objects with a similar description (for example, “wall cabinet” and “base cabinet”).
 Also, when assessing the quality of semantic segmentation, it turned out that the methods show worse results if we use a ‘_’ instead of a space in the class name.
-Object merging issues occur, for example in ConceptGraphs on ReplicaCAD, where staircases are split into individual steps. This impacts Measurement-based question answering, e.g.:
+There is also an object merging issue that affects the answers to Measurement-based questions. For example, in ReplicaCAD, both methods mark each step of the stairs as a separate staircase, e.g.:
 Question: How many staircases are present?
@@
 Scene Graph Answer: 15
-Similar errors appear in BBQ, inflating instance counts.
 VI Conclusion
-In this paper, we target a systematic evaluation of open semantic mapping algorithms, in particular their robustness in dynamic environments, e.g.
-under varying indoor lighting conditions.
-To accomplish the task we introduced a dynamically configurable and highly automated LLM/LVLM-powered benchmarking methodology.
-Our novel dataset enabled a comprehensive analysis of leading methods, such as OpenScene [peng2023openscene], BBQ [linok2024beyond], and Concept Graph [gu2024conceptgraphs], highlighting the impact of lighting variations on semantic segmentation accuracy and quality of constructed knowledge graphs.
+In this paper, we target a systematic evaluation of open semantic mapping algorithms, in particular their robustness in dynamic environments, e.g., under varying indoor lighting conditions.
+To accomplish the task, we introduced a dynamically configurable and highly automated LLM/LVLM-powered benchmarking methodology.
+Our benchmarking procedure enabled a comprehensive analysis of leading methods, such as OpenScene [3], BBQ [2], and Concept Graph [1], highlighting the impact of lighting variations on semantic segmentation accuracy and constructed knowledge graphs quality.
 Analysis of the obtained results and corner cases points directions for further research in developing more resilient algorithms capable of maintaining high performance across diverse environments.
-In our future work, we aim to expand the proposed methodology to assess a broader range of OSM methods focusing on skills crucial for autonomous navigation and general-purpose robots in dynamic environments.
-In particular, benchmarking visual grounding and scene graph evaluation skills will require adapting the suggested VQA pipeline either to generate question based on ground truth info on all instances and its positions (if available from the dataset) or to process a complete scene representation rather than analyzing it frame-by-frame.
-Verifying actionable affordances, spatial reasoning, and traversability skills will require only extending questions’ categories within the VQA pipeline, i.e.
-adapting the service prompts.
-We also did a preliminary study on incorporating synthetic dynamic objects into scenes and will continue to develop this aspect for the upcoming edition of OSMa-Bench.
-[…]
+Our solution offers a flexible framework that allows easy extension to evaluate additional capabilities of OSM approaches.
+OSMa-Bench’s adaptability allows it to generate scenes with dynamic objects, evaluate visual grounding, and assess spatial reasoning by expanding the question categories within the existing pipeline.
+We hope that our code implementation will be useful for further research.